# Verifying LLM Inference to Detect Model Weight Exfiltration

### v2

@@ B.2 Gumbel-Max FSSLEstimators (FSSL-GM)
1:Input: Claimed token $t^{*}\in V$, verifier logits $\hat{\ell}\in\mathbb{R}^{V}$, PRNG seed $\sigma$, temperature $T$, filtering parameters $(k,p)$, active subset size $s$, noise CDFs/PDF $\{F_{\varepsilon_{i}},f_{\varepsilon_{t^{*}}}\}$, number of samples $M$
@@
4:Sample Gumbel noise $G_{\sigma,i}\sim\text{Gumbel}(0,1)$ for $i=1,\ldots,V$ using seed $\sigma$
@@
10:  $A_{\sigma,i}\leftarrow\hat{\ell}_{t^{*}}-\hat{\ell}_{i}+T(G_{\sigma,t^{*}}-G_{\sigma,i})$
@@
18:   $\pi_{t^{*},j}\leftarrow\prod_{i\in S}F_{\varepsilon_{i}}\left(\max(x_{j}+A_{\sigma,i},B_{i})\right)$
@@
B.3 Naive Methods for Sampled Token Verification
@@
1:input sequence $x$, target output sequence $y$, RNG seed $s$, model $\theta$
@@
3:Initialize model RNG with seed $s$
@@
5:for $i\leftarrow 1$ to $|y|$ do
6:  $\text{next}\leftarrow\text{SampleNextToken}(\theta,\ \text{tokens}) \triangleright$ sampling deterministic given $s$
7:  if $\text{next}=y_{i}$ then
8:   $\text{diff}\leftarrow\text{diff}+1$
9:  end if
10:  $\text{tokens}\leftarrow\text{tokens}\circ y_{i} \triangleright$ teacher forcing: append ground-truth token
@@
Algorithm 6 Sampled token verification via prefill
1:input sequence $x$, target output $y=(y_{1},\dots,y_{T})$, RNG seed $s$, model $\theta$
2:Initialize RNG with seed $s$
[…]
6:for $i\leftarrow 1$ to $T$ do $\triangleright$ predict $y_{i}$ from context $x\circ y_{<i}$
7:  $p_{i}\leftarrow\text{Softmax}\!\left(\ell_{|x|+i-1}\right)$
8:  $\hat{y}_{i}\leftarrow\text{SampleFrom}(p_{i})$ $\triangleright$ sampling draws consume RNG in fixed order
9:  if $\hat{y}_{i}=y_{i}$ then
10:   $\text{diff}\leftarrow\text{diff}+1$
11:  end if
12:end for
[…]
14:return $\text{diff},\ \text{score}$
B.4 Rank-Aware FSSL-GM Scoring
@@
Appendix C Mechanism Design security solutions: Random Sampling and the Panopticon
C.1 Mechanism Design
@@
C.2 Derivation of Required Sampling Rate
@@
Appendix D Trust assumptions for Sampled Output Verification
D.1 Assumption 1: The Verification Server is trusted to correctly execute its code
In our proposed solution, it is important that the verification server correct executes its code. When the verification code itself is untrusted or run in an untrusted environment, an adversary could potentially compromise and circumvent the policy enforcement.
@@ D.1 Assumption 1: The Verification Server is trusted to correctly execute its code
D.2 Assumption 2: All messages passed between users and the Inference Server are correctly logged and securely stored
@@
D.3 Assumption 3: The attacker can neither predict nor alter which logs the Verification Server will sample to verify
It is critical that an adversary attempting to exfiltrate weights does not know which outputs are going to be audited. Otherwise, an adversary may simply exfiltrate weights through the messages not-audited. This assumption is critical for this scheme to work; however, many schemes for random sampling exist. A simple proposal like random sampling within a trusted-execution environment is almost certainly sufficient; but in the event that it isn’t this has been studied in great depth in blockchain literature. We discuss how to do this in detail in Section D.6.
D.4 Assumption 4: The Verification Server itself does not pose a risk of model exfiltration
@@
D.5 Assumption 5: The randomness used for inference is trusted and non-malleable
If the adversary can manipulate randomness, they can transmit information by selecting seeds/samples while leaving per-step likelihoods within normal ranges. In this regime the verifier’s guarantees degrade sharply: many exfiltration strategies become indistinguishable from benign stochasticity, and a malicious provider could even watermark or backdoor outputs via RNG tampering while remaining white-box undetectable [goldwasser2024plantingundetectablebackdoorsmachine, pseudo_random_error_correcting].
@@ D.5 Assumption 5: The randomness used for inference is trusted and non-malleable
If an adversary can steer the randomness (e.g., select seeds, bias PRNG state, or grind over retries), the sampler itself becomes a covert channel: the adversary can encode arbitrary payloads by “choosing the random samples.” Closely related phenomena are known in training-time settings, where tampering with algorithmic randomness enables undetectable backdoors that are indistinguishable to any efficient observer [goldwasser2024plantingundetectablebackdoorsmachine]. Analogously, inference-time coin-fixing enables steganographic signaling without changing logits in a way a verifier could reliably flag (see also pseudo-random watermarking constructions [pseudo_random_error_correcting]).
@@
D.6 Message-Sampling Secure Seed Expansion
@@
    Expand the seed adaptively, based on tokens chosen previously. So, computing the randomness for the next token as $H(K,\sigma,x)$, where $x$ is the prompt and tokens computed thus far (used in Theorems 4.2 and 4.3).
@@
D.7 Plausibility of Necessary Assumptions
Theorem 4.1, and to a lesser extent Theorems 4.2 and 4.3, give the adversary exfiltrating data a significant amount of power. For example, we assume both the sender and receiver have access to the model distribution for the given prompt and sometimes even the distribution of honest responses accounting for valid non-determinism.
@@ D.7 Plausibility of Necessary Assumptions
Appendix E How Specific Implementations may Affect Verification
The exact inference stack used dramatically affects the requirements on the verification server. In this work we formulate sampling from an LLM as a two step process: (i) generate a probability distribution, and (ii) sample a single token from that probability distribution. The majority of this work is showing how to quantify the non-determinism that may arise from slightly different probability distribution generating processes. However, for this to work, this is predicated on a precise and easy-to-reproduce method of sampling from the probability distribution — in this work we show how to do this for the Inverse Probability Transform method of sampling and the Gumbel-Max trick.
Modern inference stacks are generally more complicated than either of these two methods, employing tricks like speculative decoding, and this needs to be shared across the inference server and verification server, which we discuss in Section E.1.
Another noteworthy consideration is how random samplers update their state, and synchronizing this across the inference and verification servers. For example, in some implementations (like Ollama [ollama_docs]), if you filter the probability distribution with top-p and the token array ends up with only one entry, the inference server will not sample at all, and use a greedy decoding step. This means that the RNG state does not advance for the inference server, but may for the verification server. However, this issue does not arise in other common implementations like vLLM [vllm].
E.1 Speculative Decoding
Speculative decoding [leviathan2023fastinferencetransformersspeculative, chen2023acceleratinglargelanguagemodel, zhang2025learningharmonizedrepresentationsspeculative, li2025eagle3scalinginferenceacceleration] has emerged as a widely-adopted technique to accelerate LLM inference by using a smaller “draft” model to propose candidate tokens, which are then verified in parallel by the larger “target” model. This approach can achieve substantial speedups (often $1.5{-}2.5\times$) while preserving the exact output distribution of standard autoregressive sampling from the target model.
@@ E.1 Speculative Decoding
    Production systems exhibit heterogeneity in their acceleration strategies—while speculative decoding is common, implementations vary significantly across frameworks and deployments.
@@
In speculative decoding introduced by [leviathan2023fastinferencetransformersspeculative, chen2023acceleratinglargelanguagemodel], a draft model generates $K$ candidate tokens, which the target model then verifies in parallel. For each candidate, the target model either accepts it, rejects and resamples from an adjusted distribution, or (if all candidates are accepted) samples one additional bonus token. This process maintains exact distributional equivalence to standard autoregressive sampling from the target model alone.
Different speculative decoding algorithms use different methods for accepting tokens from the draft model. For example, in the original work proposing speculative decoding [leviathan2023fastinferencetransformersspeculative] the verifier only accepts a token if it is the exact token that the verifier model would have generated in that position. However, in other methods a draft token is accepted not only if it is the token the verifier would have sampled, but alternatively if it is a sufficiently likely token as well [li2025eagle3scalinginferenceacceleration].
@@
Appendix F Response to Detecting Exfiltration Attempts
This work focuses primarily on the detection of potential model-weight exfiltration attempts rather than on post-detection remediation. In particular, we formalize the task of determining whether a batch of model outputs contains covertly embedded information. Following the literature on steganographic detection [batch_stego_2007, batch_stego_real_world_2012], we envision verification pipelines adhering to a general three-step structure:
@@ Appendix F Response to Detecting Exfiltration Attempts
F.1 Privacy-preserving data analysis.
@@
Appendix G Gumbel-Likelihood Score Statistics
@@
G.1 Score Distribution for a Single Token Position
Figure 11 shows the distribution of FSSL-GM scores across the entire vocabulary for a single token position during generation. For each token $i$ in the vocabulary, we compute $\mathrm{FSSL-GM}_{t}(i)$ using the Gumbel-Max likelihood estimation procedure described in Algorithm 2.
@@ G.1 Score Distribution for a Single Token Position
G.2 Distribution of FSSL-GM Scores for Actually-Generated Tokens
@@
G.3 Threshold Selection and FPR-Information Trade-off
@@
Appendix H Gumbel-Likelihood Score Exfiltration Results
@@
Figure 14: Multi-model two-tier classifier Pareto frontier (envelope across all rank thresholds). This plot displays the optimal security-vs-usability trade-off for detecting steganographic data exfiltration across multiple LLM architectures. Each curve represents a different model, showing the envelope across all tested rank thresholds—i.e., for each FPR level, it shows the minimum extractable information achievable by choosing the optimal rank threshold. The Pareto frontier shows the best achievable trade-off between (x-axis) False Positive Rate (%) - fraction of legitimate tokens incorrectly flagged as “dangerous” and (y-axis) Extractable Information (%) - how much data an attacker can still exfiltrate while evading detection. Models tested: Llama-3.1-8B, Llama-3.2-3B, Qwen3-30B-A3B, and Qwen1.5-MoE-A2.7B.
H.1 Logit-Rank Ablations
@@
H.2 Context Length Ablations
To understand how input context length affects the FSSL-GM verification system’s performance, we conducted ablation studies using prompts of varying lengths. We generated prompts with 4,000, 8,000, and 12,000 tokens using the LongBench v2 dataset [bai2025longbenchv2deeperunderstanding], which provides realistic long-context tasks for multitask benchmarks. We create generations of max-length 500 tokens for 100 prompts, given 100 FSSL-GMthresholds, and $\sigma\in(0.001,0.01,0.1,1.)$. For each context length and model combination, we measured the trade-off between false positive rate and extractable information under the same experimental setup as Section 6.
@@ H.2 Context Length Ablations
We draw this parallel as a potential future direction for developing future exfiltration mitigation. Current watermarking schemes are built for different infrastructure and generally designed to be robust to perturbations [ARXIV:KGWKMG23, COLT:ChrGunZam24, SP:CHS25]. However, it is an interesting possibility for future work to adapt techniques from this literature to this setting.
As a philosophically interesting point, we observe that one can view the problem of verifying outputs as purposefully producing an extremely brittle watermarking scheme — namely, one that any text generated by the model is seen as “watermarked” and any deviation from that text is no longer seen as watermarked.
It is valuable to note that our verification solution is entirely compatible with current watermarking schemes too. So, an AI system which is implementing watermarking already can implement this on top of it, with no changes to the system — though, if the watermarking scheme involves a private key, that private key would need to be reproduced in the verification server.
Appendix I Non-Exhaustive Taxonomy of Exfiltration Vectors
@@
I.1 Main Message Channel
@@ I.2 Developer Channels
I.3 Operational Security Channel
@@ I.5 Miscellaneous
Appendix J Recommended (High-level) Defenses
@@
J.1 More Concrete Recommendations
@@
Figure 26: Egress-limiting schematic (illustrative): controls on outbound channels reduce bandwidth and increase observability for exfiltration attempts. Egress limiting as a form of strong guarantees on exfiltration is employed by the company Anthropic [anthropic2025asl3].

### v3

@@ C.2 Gumbel-Max FSSLEstimators (FSSL-GM)
1:Input: Claimed token $t^{*}\in\mathcal{T}$, verifier logits $\hat{\ell}\in\mathbb{R}^{|\mathcal{T}|}$, Seed $\mathsf{seed}$, temperature $T$, filtering parameters $(k,p)$, active subset size $s$, noise CDFs/PDF $\{F_{\varepsilon_{i}},f_{\varepsilon_{t^{*}}}\}$, number of samples $M$
@@
4:Sample Gumbel noise $G_{\mathsf{seed},i}\sim\text{Gumbel}(0,1;\mathsf{H}(\mathsf{seed}\|i))$ for $i=1,\ldots,|\mathcal{T}|$
@@
10:  $A_{\mathsf{seed},i}\leftarrow\hat{\ell}_{t^{*}}-\hat{\ell}_{i}+T(G_{\mathsf{seed},t^{*}}-G_{\mathsf{seed},i})$
@@
18:   $\pi_{t^{*},j}\leftarrow\prod_{i\in S}F_{\varepsilon_{i}}\left(\max(x_{j}+A_{\mathsf{seed},i},B_{i})\right)$
@@
C.3 Naive Methods for Sampled Token Verification
@@
1:input sequence $\mathcal{H}$, target output sequence $y$, seed $\mathsf{seed}$, model $\theta$
@@
3:for $i\leftarrow 1$ to $|y|$ do
4:  $\text{next}\leftarrow\text{SampleNextToken}(\theta,\mathcal{H};\mathsf{H}(\mathsf{seed}\|i))$
5:  if $\text{next}=y_{i}$ then
6:   $\text{diff}\leftarrow\text{diff}+1$
7:  end if
8:  $\mathcal{H}\leftarrow\mathcal{H}\|y_{i} \triangleright$ teacher forcing: append ground-truth token
9:end for
@@
11:return $\text{diff},\ \text{score}$
Algorithm 6 Sampled token verification via prefill
1:input sequence $x$, target output $y=(y_{1},\dots,y_{T})$, seed $\mathsf{seed}$, model $\theta$
[…]
5:for $i\leftarrow 1$ to $T$ do $\triangleright$ predict $y_{i}$ from context $x\circ y_{<i}$
6:  $p_{i}\leftarrow\text{Softmax}\!\left(\ell_{|x|+i-1}\right)$
7:  $\hat{y}_{i}\leftarrow\text{SampleFrom}(p_{i};\mathsf{H}(\mathsf{seed}\|i))$
8:  if $\hat{y}_{i}=y_{i}$ then
9:   $\text{diff}\leftarrow\text{diff}+1$
10:  end if
@@
C.4 Rank-Aware FSSL-GM Scoring
@@
Appendix D Mechanism Design security solutions: Random Sampling and the Panopticon
D.1 Mechanism Design
@@
D.2 Derivation of Required Sampling Rate
@@
Appendix E Trust assumptions for Sampled Output Verification
E.1 Assumption 1: The Verification Server is trusted to correctly execute its code
In our proposed solution, it is important that the verification server correctly executes its code. When the verification code itself is untrusted or run in an untrusted environment, an adversary could potentially compromise and circumvent the policy enforcement.
@@ E.1 Assumption 1: The Verification Server is trusted to correctly execute its code
E.2 Assumption 2: All messages passed between users and the Inference Server are correctly logged and securely stored
@@
E.3 Assumption 3: The attacker can neither predict nor alter which logs the Verification Server will sample to verify
It is critical that an adversary attempting to exfiltrate weights does not know which outputs are going to be audited. Otherwise, an adversary may simply exfiltrate weights through the messages not-audited. This assumption is critical for this scheme to work; however, many schemes for random sampling exist. A simple proposal like random sampling within a trusted-execution environment is almost certainly sufficient; but in the event that it isn’t this has been studied in great depth in blockchain literature. We discuss how to do this in detail in Section E.6.
E.4 Assumption 4: The Verification Server itself does not pose a risk of model exfiltration
@@
E.5 Assumption 5: The randomness used for inference is trusted and non-malleable
If the adversary can manipulate randomness, they can transmit information by selecting seeds/samples while leaving per-step likelihoods within normal ranges. In this regime the verifier’s guarantees degrade sharply: many exfiltration strategies become indistinguishable from benign stochasticity, and a malicious provider could even watermark or backdoor outputs via RNG tampering while remaining white-box undetectable [GKV+24, CG24].
@@ E.5 Assumption 5: The randomness used for inference is trusted and non-malleable
If an adversary can steer the randomness (e.g., select seeds, bias PRNG state, or grind over retries), the sampler itself becomes a covert channel: the adversary can encode arbitrary payloads by “choosing the random samples.” Closely related phenomena are known in training-time settings, where tampering with algorithmic randomness enables undetectable backdoors that are indistinguishable to any efficient observer [GKV+24]. Analogously, inference-time coin-fixing enables steganographic signaling without changing logits in a way a verifier could reliably flag (see also pseudo-random watermarking constructions [CG24]).
@@
E.6 Message-Sampling Secure Seed Expansion
@@
    Expand the seed adaptively, based on tokens chosen previously. So, computing the randomness for the next token as $H(K,\sigma,x)$, where $x$ is the prompt and tokens computed thus far (used in Theorems 5.2 and 5.3).
@@
E.7 Plausibility of Necessary Assumptions
Theorem 5.1, and to a lesser extent Theorems 5.2 and 5.3, give the adversary exfiltrating data a significant amount of power. For example, we assume both the sender and receiver have access to the model distribution for the given prompt and sometimes even the distribution of honest responses accounting for valid non-determinism.
@@ E.7 Plausibility of Necessary Assumptions
Appendix F How Specific Implementations may Affect Verification
The exact inference stack used dramatically affects the requirements on the verification server. In this work we formulate sampling from an LLM as a two step process: (i) generate a probability distribution, and (ii) sample a single token from that probability distribution. The majority of this work is showing how to quantify the non-determinism that may arise from slightly different probability distribution generating processes. However, for this to work, this is predicated on a precise and easy-to-reproduce method of sampling from the probability distribution. In this work we show how to do this for the Inverse Probability Transform method of sampling and the Gumbel-Max trick.
Modern inference stacks are generally more complicated than either of these two methods, employing tricks like speculative decoding, and this needs to be shared across the inference server and verification server, which we discuss in Section F.1.
Another noteworthy consideration is how random samplers update their state, and synchronizing this across the inference and verification servers. For example, in some implementations (like Ollama [32]), if you filter the probability distribution with top-p and the token array ends up with only one entry, the inference server will not sample at all, and use a greedy decoding step. This means that the RNG state does not advance for the inference server, but may for the verification server. However, this issue does not arise in other common implementations like vLLM [KLZ+23].
F.1 Speculative Decoding
Speculative decoding [LKM23, CBI+23, ZWH+25, LWZ+25] has emerged as a widely-adopted technique to accelerate LLM inference by using a smaller “draft” model to propose candidate tokens, which are then verified in parallel by the larger “target” model. This approach can achieve substantial speedups (often $1.5{-}2.5\times$) while preserving the exact output distribution of standard autoregressive sampling from the target model.
@@ F.1 Speculative Decoding
    Production systems exhibit heterogeneity in their acceleration strategies: while speculative decoding is common, implementations vary significantly across frameworks and deployments.
@@
In speculative decoding introduced by [LKM23, CBI+23], a draft model generates $K$ candidate tokens, which the target model then verifies in parallel. For each candidate, the target model either accepts it, rejects and resamples from an adjusted distribution, or (if all candidates are accepted) samples one additional bonus token. This process maintains exact distributional equivalence to standard autoregressive sampling from the target model alone.
Different speculative decoding algorithms use different methods for accepting tokens from the draft model. For example, in the original work proposing speculative decoding [LKM23] the verifier only accepts a token if it is the exact token that the verifier model would have generated in that position. However, in other methods a draft token is accepted not only if it is the token the verifier would have sampled, but alternatively if it is a sufficiently likely token as well [LWZ+25].
@@
Appendix G Response to Detecting Exfiltration Attempts
This work focuses primarily on the detection of potential model-weight exfiltration attempts rather than on post-detection remediation. In particular, we formalize the task of determining whether a batch of model outputs contains covertly embedded information. Following the literature on steganographic detection [KER07, KP12], we envision verification pipelines adhering to a general three-step structure:
@@ Appendix G Response to Detecting Exfiltration Attempts
G.1 Privacy-preserving data analysis.
@@
Appendix H FSSL-GM Score Statistics
@@
H.1 Score Distribution for a Single Token Position
Figure 11 shows the distribution of FSSL-GM scores across the entire vocabulary for a single token position during generation. For each token $i$ in the vocabulary, we compute $\mathrm{FSSL-GM}_{t}(i)$ using the Token-DiFR estimation procedure described in Algorithm 4.
@@ H.1 Score Distribution for a Single Token Position
H.2 Distribution of FSSL-GM Scores for Actually-Generated Tokens
@@
H.3 Threshold Selection and FPR-Information Trade-off
@@
Appendix I Gumbel-Likelihood Score Exfiltration Results
@@
Figure 14: Multi-model two-tier classifier Pareto frontier (envelope across all rank thresholds). This plot displays the optimal security-vs-usability trade-off for detecting steganographic data exfiltration across multiple LLM architectures. Each curve represents a different model, showing the envelope across all tested rank thresholds, i.e., for each FPR level, it shows the minimum extractable information achievable by choosing the optimal rank threshold. The Pareto frontier shows the best achievable trade-off between (x-axis) False Positive Rate (%) - fraction of legitimate tokens incorrectly flagged as “dangerous” and (y-axis) Extractable Information (%) - how much data an attacker can still exfiltrate while evading detection. Models tested: Llama-3.1-8B, Llama-3.2-3B, Qwen3-30B-A3B, and Qwen1.5-MoE-A2.7B.
I.1 Logit-Rank Ablations
@@
I.2 Context Length Ablations
To understand how input context length affects the FSSL-GM verification system’s performance, we conducted ablation studies using prompts of varying lengths. We generated prompts with 4,000, 8,000, and 12,000 tokens using the LongBench v2 dataset [BTZ+25], which provides realistic long-context tasks for multitask benchmarks. We create generations of max-length 500 tokens for 100 prompts, given 100 FSSL-GMthresholds, and $\sigma\in(0.001,0.01,0.1,1.)$. For each context length and model combination, we measured the trade-off between false positive rate and extractable information under the same experimental setup as Section 7.
@@ I.2 Context Length Ablations
We draw this parallel as a potential future direction for developing future exfiltration mitigation. Current watermarking schemes are built for different infrastructure and generally designed to be robust to perturbations [KGW+23a, CGZ24, CHS25]. However, it is an interesting possibility for future work to adapt techniques from this literature to this setting.
As a philosophically interesting point, we observe that one can view the problem of verifying outputs as purposefully producing an extremely brittle watermarking scheme, namely, one that any text generated by the model is seen as “watermarked” and any deviation from that text is no longer seen as watermarked.
It is valuable to note that our verification solution is entirely compatible with current watermarking schemes too. So, an AI system which is implementing watermarking already can implement this on top of it, with no changes to the system, though, if the watermarking scheme involves a private key, that private key would need to be reproduced in the verification server.
Appendix J Empirical Base Rates of non-determinism
Empirical Characterization of Non-Determinism.
To quantify the inherent non-determinism in language model outputs, we generate 1,000 responses to the same prompt using identical random seeds and construct a token-level prefix tree to visualize divergence patterns (Figure 26(b)).
Each node in the tree represents a unique token prefix shared by one or more responses. Nodes are connected to their continuations, forming branches that capture all distinct response trajectories. When responses sharing a prefix diverge into multiple next-token choices, we mark that position as a fork point and record its fork factor—the number of unique next tokens (i.e., the number of children). Nodes with a single child correspond to deterministic continuations, whereas fork points indicate positions where the model produced different tokens despite identical seeds.
Our primary experiments are conducted with OpenAI 4o-mini, which exhibits a moderate but measurable level of seeded non-determinism. Across 1,000 runs, we observe a “mostly deterministic with critical decision points” pattern: approximately 95% of nodes have exactly one child, 5% represent binary forks, and only 0.2% involve higher-order splits. This indicates that even under fixed seeds, small implementation-level or system-level variations can lead to isolated yet systematic divergences in generation trajectories.
(a) Distribution of fork factors for a prompt.
(b) Prefix tree visualization of 1,000 responses to the same prompt under identical random seeds (first 100 tokens shown). Red nodes mark fork points (multiple children), with size proportional to fork factor.
Figure 26: Empirical characterization of seeded non-determinism for a single prompt using OpenAI 4o-mini. (26(a)) The histogram quantifies the overall frequency of fork points, while (26(b)) the prefix tree visualizes how and where these divergences occur in token space.
Observations.
Empirically, these results demonstrate that valid non-determinism in language model outputs is highly structured and non-uniform. Rather than exhibiting uniform stochasticity across tokens, variability concentrates at specific decision points. Most tokens are effectively deterministic—reproducing the same output across all runs—while a small minority exhibit localized randomness, occasionally branching into several (up to five or more) distinct continuations.
Interestingly, when we replicate this experiment locally using vLLM with Llama-3.1-8B, the generation process becomes dramatically more deterministic (unsurprisingly, as both system load and numerical environment are tightly controlled). In this setting, 99.8% of nodes have a single child, and only 0.2% exhibit any branching at all. This contrast reinforces that much of the observed non-determinism in hosted models such as 4o-mini arises not from the model’s sampling rule itself, but from small, environment-level or distributed-inference sources of variability.
Appendix K Non-Exhaustive Taxonomy of Exfiltration Vectors
@@
K.1 Main Message Channel
@@
K.3 Operational Security Channel
@@
Appendix L Recommended (High-level) Defenses
@@
L.1 More Concrete Recommendations
@@
Figure 27: Egress-limiting schematic (illustrative): controls on outbound channels reduce bandwidth and increase observability for exfiltration attempts. Egress limiting as a form of strong guarantees on exfiltration is employed by the company Anthropic [ANT25].